Solve the equation by using the Square Root Property.
13.) 16x^2 - 48x + 36 = 49

Answer by checkley77(12844)

16x^2 - 48x + 36 = 49
(4x-6)^2=7^2 taking the square roots we get:
4x-6=7
4x=7+6
4x=13
x=13/4